### v1

@@ 4.1 Visual inspection
The dust-to-metal mass ratio ${\rm DTM}=\rho_{\rm D}/\rho_{\rm Z}$ and the dust-to-gas mass ratio ${\rm DTG}=\rho_{\rm D}/\rho$, where $\rho_{\rm D}$ is the total dust mass density, and $\rho_{\rm Z}$ is the total metal (dust+gas) density, show significant spatial variations in the ISM. The DTM is higher in regions of higher densities due to the increase in accretion rates of refractory elements on dust with gas density, and to the higher destruction rates by thermal sputtering at higher temperatures. Dense gas also corresponds to regions where the mass fraction of small grains $f_{\rm S}$ is lower and the mass fraction of carbonaceous grains $f_{\rm C}$ is higher compared to the more diffuse ISM.
@@ 4.2 Dust-to-metal ratio
The evolution over time of the DTM for the fiducial G10LG galaxy is shown in Fig. 3. This quantity is computed for all the gas and dust contained in a cylinder of $r=4\,\rm kpc$ cylindrical radius centered on the box and $h/2=200\,\rm pc$ above and below the disc plane. The DTM rises steeply in less than $50\,\rm Myr$ to its approximately steady-state value at nearly ${\rm DTM}=0.4$ comparable to the canonical ratio in present-day galaxies with solar-like metallicities (e.g. Rémy-Ruyer et al. 2014; De Vis et al. 2019) or in the Milky Way (e.g. Jenkins 2009). The DTM is decomposed into the four simulated dust bins, i.e. into small (dotted) and large (dashed) grains, and into carbonaceous (blue) and silicate (red) grains. The bulk of the DTM is from large grains ($80\,\%$) and from silicates ($66\,\%$), and in particular from large silicate grains ($53\,\%$).
@@ 4.3 Dust size distribution
The top panel of Fig. 5 shows the reconstructed size distribution of grains from our four populations with size $a_{j}$ ($j$ standing for S or L), and chemical composition $i$ (for carbonaceous or silicate grains), assuming that each population is sampled with a modified log-normal PDF (as in Hirashita 2015 and Hou et al. 2017) of
@@
where $\sigma_{i,j}=0.75$ is the standard deviation of the log-normal part, and $C_{i,j}$ is the normalisation of the distribution. The normalisation is given by:
@@
where $\mu=1.22$ is the mean molecular weight of the gas (assuming, for simplicity, that the dusty gas mainly contributing to the extinction curve is fully neutral). The size distribution of grains measured by $\mathcal{D}_{i,j}=\sum_{\Omega}D_{i,j}/\sum_{\Omega}M_{\rm g}$, over a control volume $\Omega$ that is a cylinder of $r=4\,\rm kpc$ radius and $h/2=0.2\,\rm kpc$ semi-height around the center of the galaxy, is in good agreement with the expected MRN distribution of grains in the Milky Way. It shows that the size distribution of carbonaceous and silicate grains are extremely similar, and that the amount of carbonaceous grains is lower than that of silicate grains (as expected from the decomposition of DTM in Fig. 3).
@@ 4.4 Extinction curve
where $N_{\rm H}$ is the hydrogen column number density, and $Q_{{\rm ext},i}(a,\lambda)$ is the extinction coefficient that is a function of grain size $a$, wavelength $\lambda$ and grain composition $i$. The size distribution is obtained from the log-normal distribution [^7] Note that the extinction curve is not sensitive to the exact spread of the log-normal reconstruction for the small bin size, but is very sensitive to the one for the large bin (and in particular to that of carbonaceous grains), which should have a sharp cut-off above $a=0.1\,\mu\rm m$ (Weingartner & Draine 2001a). of equation (21). $Q_{{\rm ext},i}(a,\lambda)$ is obtained by using the Mie theory (Bohren & Huffman 1983) with the optical constants for carbonaceous and silicate grains from Weingartner & Draine (2001a). Bottom panel of Fig. 5 shows the extinction curve $A_{\lambda}$ normalised by the extinction $A_{\rm V}$ in the V-band at $\lambda_{\rm V}=0.55\,\rm\mu m$, obtained for the fiducial G10LG simulation at time $t=100$, 200, 300, and $400\,\rm Myr$ for all the dust in the same region (within a cylinder of $r=4\,\rm kpc$ and $h/2=0.2\,\rm kpc$) as used previously for measuring the dust size distribution (top panel of Fig. 5). We also show the scatter of the extinction curve at $t=400\,\rm Myr$ by sampling the corresponding extinction curves for each individual cells within the same region. The extinction curve from the Milky Way analogue G10LG shows an excellent agreement with that observed for the Milky Way (Pei 1992): it shows a similar slope at short wavelengths and the characteristic bump feature at $\lambda=2175\,\rm\AA$ with similar strength. This shape is clearly distinct from the extinction curves from Magellanic Clouds. As expected, the $2175\,\rm\AA$ bump is obtained from the small carbonaceous grains (blue dotted line) due to their optical properties, while the UV-to-optical slope is the result of the small silicate grains (red dotted line). It follows immediately that to obtain steeper slopes and a shallower $2175\,\rm\AA$ bump (as in the Magellanic Clouds extinction curves), the fraction of small carbonaceous grains must be reduced, and the fraction of small silicate grains must be enhanced. This will be investigated further in section 5.
@@
Diversity in the extinction curves is also a significant feature of Milky Way extinction curves with typical relative variations across lines-of-sights of $A_{\lambda}/A_{\rm V}$ of $\sim 20\,\%$ at $\lambda^{-1}=8\,\mu\rm m^{-1}$ (Cardelli et al. 1989; Fitzpatrick & Massa 2007), as can be seen in Fig. 5. Although, extreme care must be taken when comparing directly with observational data (line-of-sight effects), we investigate what are the different extinction curves as a function of the various gas phases of the ISM to get a better indication of what are variations of the dust properties across them. In Fig. 6, we show the extinction curves at $t=400\,\rm Myr$ in G10LG, as relative variations of $\delta(A_{\lambda}/A_{\rm V})|_{n}=(A_{\lambda}/A_{\rm V})|_{n}/(A_{\lambda}/A_{\rm V})-1$ at a given $n$ with respect to the global extinction curve $A_{\lambda}/A_{\rm V}$ (as shown in Fig. 5), for different mean gas densities $n\sim 10^{-2},10^{-1},1,10$, and $10^{2}\,\rm H\,cm^{-3}$ (sampled in narrow bins of ±$\pm 0.1\,\rm dex$ around the mean) collected within the same cylinder of $r=4\,\rm kpc$ and $h/2=0.2\,\rm kpc$. Gas at densities of $n\sim 10$ and $10^{2}\,\rm H\,cm^{-3}$ exhibits extinction curves close to the global value. The variation of extinction curves is not monotonic with gas density. The extremely low value of gas density $n\sim 10^{-2}\,\rm H\,cm^{-3}$ corresponds to the larger decrement in the extinction curves: lower bump strength and shallower UV-to-optical slope. This is understood as the consequence of small grain destruction by SN explosions and thermal sputtering. At intermediate densities $n\sim 10^{-1}$ and $1\,\rm H\,cm^{-3}$, the extinction curves approach their global values, and stand above the global expectation at $1\,\rm H\,cm^{-3}$ as a signature of shattering. Shattering transfers mass from large grains to small grains preferentially in the diffuse ISM, reinforcing both the bump strength and the UV slope. The highest gas densities are the closest to the global value as this is where most of the dust mass is concentrated ($50\,\%$ of the total dust mass at $n\sim 10\,\rm H\,cm^{-3}$, and $30\,\%$ at $n\sim 10^{2}\,\rm H\,cm^{-3}$). They have lower extinction curves relative to the lower ISM gas density of $n\sim 1\,\rm H\,cm^{-3}$ due to the efficient coagulation of small grains into large grains. Finally the different sampled densities all have large scatter of the order of the mean values, which is indicative of the intrinsic diversity of extinction curves at a given gas density.
[…]
@@ 4.5 Depletion factors
The presence of dust in the ISM can be characterised by the depletion of the corresponding atomic elements in the gas phase $f_{\rm dep}(X)=1-M_{\rm D}(X)/M_{\rm Z}(X)$, where $M_{\rm D}(X)$ and $M_{\rm Z}(X)$ are respectively the dust mass and the total (dust+gas) mass of the element $X$. We show in Fig. 7 the depletion factor of the various elements as a function of the neutral gas density $n_{\rm neutral}$ (assuming that gas with temperature below $T\leq 10^{4}\,\rm K$ is neutral and fully ionised otherwise) compared to observations. The observational relations are shown with the raw value of Jenkins (2009) as the black dashed lines, and with the recommended renormalisation of their data in black solid line by Zhukovska et al. 2016 (a correction also adopted in Choban et al. 2022) that compensates for the fact that the observed densities are under-estimated since they are averaged along the line-of-sight. We also show the fit provided by Richings et al. (2022) of the data from De Cia et al. (2016), which extends the Jenkins (2009) data to low density lines-of-sight (with or without the rescaling in solid and dashed red lines respectively). For C depletion, we follow Choban et al. (2022), and we reduce the depletion values of Jenkins (2009) by a factor 2 (in that case, we do not rescale in density because the relation is sufficiently shallow) to account for the apparent over-estimate of the C gas-phase abundance from weak compared to strong CII lines (Sofia et al. 2011; Parvathi et al. 2012). We also show the data points from Jenkins (2009) (rescaled by a factor of 2) and the data points from Parvathi et al. (2012) to underline the large uncertainties associated to the estimate of the depletion of C.
[…]
@@
For C depletion, our resulting values are well within the range of observations although the observational data exhibits a large scatter. We also highlight the role of CO formation at high gas densities in limiting the depletion of C elements by grains (see also Choban et al. 2022). If we do not account for CO formation in the limitation of carbonaceous dust growth from ISM accretion, i.e., as for silicate grains, carbonaceous grain growth is only limited by ice mantle coating starting above $n=10^{4}\,\rm H\,cm^{-3}$ (G10LG_NCO simulation), then carbonaceous grains accrete much more efficiently from the ISM, leading to too much depletion of C. Therefore, CO formation indirectly allows for more moderate values of $f_{\rm dep}({\rm C})$ in better agreement with the data.
@@ 4.6 Which process dominates the dust growth?
Figure 9: The DTM (top left), the DTG (top right) the fraction of small grains $f_{\rm S}$ (bottom left) and the fraction of carbonaceous grains $f_{\rm C}$ (bottom right) as a function of density $n$ and temperature $T$ for the G10LG galaxy at time $t=400\,\rm Myr$. Black contours are for the mass distribution of gas.
Figure 10: Top panel: The dust-to-metal ratio (solid lines, left axis) and the gas metallicity (dashed lines, right axis) as a function of time for the 4 different initial metallicities of the G10LG galaxy: $0.1$, $0.3$, $1$, and $2\,\rm Z_{\odot}$ (from bottom to top). Bottom panel: the fraction of carbonaceous grains (solid lines) and of small grains (dot-dashed lines). The DTM increases with metallicity, the fraction of carbonaceous grains increases with metallicity and the fraction of small grains decreases with metallicity.
Figure 11: Comparison of extinction curves in the G10LG galaxy at time $t=400\,\rm Myr$ for different initial metallicities ($Z_{\rm g,0}=0.1$, $0.3$, $1$, and $2\,\rm Z_{\odot}$ for G10LG_VLZ, G10LG_LZ, G10LG, and G10LG_HZ respectively). The top panel shows the extinction curves, and the bottom panel shows the relative variation of the extinction curve with respect to the fiducial simulation G10LG. Galaxies with lower metallicities produce steeper UV-to-optical slopes due lower accretion efficiencies in the dense gas where coagulation proceeds, hence, the fraction of small grains is larger.
@@
Figure 13: Comparison of extinction curves in the G10LG galaxy at time $t=400\,\rm Myr$ for variations in coagulation (red), shattering rates (blue), and in SN dust destruction efficiency for small grains (cyan). An increase (decrease) in the corresponding rate is depicted by a dashed (respectively solid) line. The triple-dot-dashed orange line stands for an increased in both the shattering and the coagulation rates. Top panel shows the extinction curves, and the bottom panel shows the relative variation of the extinction curve with respect to the fiducial simulation.
@@
The evolution of the dust mass in the simulated galaxies obeys different growth and destruction processes. Figure 8 shows the different dust mass variation rates for different processes involved in the dust evolution of the fiducial G10LG run: accretion, thermal sputtering, SN shock destruction, dust released by stars through stellar winds in SN ejecta, astration, and coagulation and shattering. All rates are obtained within the whole simulation box, and curves are smoothed over a 5 Myr timescale for sake of readability. Apart from the first initial peak at a few $10\,\rm Myr$ after the start of the simulation, all curves show little variation with time due to the low SFR and no significant evolution in gas-phase metallicity (which could affect the accretion time since $t_{\rm acc}\propto Z^{-1}$). Accretion of the available refractory elements from the ISM onto the small grains is by far the dominant process of dust mass growth compared to dust released in the ejecta of SNe or of stellar winds, or by accretion onto large grains. Destruction by SNe is a factor of 3 below the mass growth by accretion and is the dominant mechanism for dust mass removal. Thermal sputtering, which is the mechanism by which the dust is destroyed by pockets of hot ($T>10^{6}\,\rm K$) and diffuse ($n<0.01\,\rm H\,cm^{-3}$) gas, is a factor 2 below the destruction rate by SNe. Finally destruction of the dust by astration is lower by a factor 4 compared to destruction by SNe. Interestingly, the thermal sputtering removes more mass on large grains than on small grains, and SNII destruction removes an equal mass of small and large grains, while their physical timescale for dust destruction is an increasing function of the grain size. Hence, one would naively expect the sputtering to remove more dust mass from small grains. However, since the dust mass is 4 times larger for large grains than small grains (see Fig. 3), this effect is compensated by the available reservoir of dust mass in each bin of grain size.
@@
4.7 Dust grain distribution in the multiphase ISM

### v2

@@ 4.1 Visual inspection
The dust-to-metal mass ratio ${\rm DTM}=\rho_{\rm D}/\rho_{\rm Z}$ and the dust-to-gas mass ratio ${\rm DTG}=\rho_{\rm D}/\rho$, where $\rho_{\rm D}$ is the total dust mass density and $\rho_{\rm Z}$ is the total metal (dust+gas) density, show significant spatial variations in the ISM. The DTM is higher in regions of higher densities due to the increase in accretion rates of refractory elements on dust with gas density, and to the higher destruction rates by thermal sputtering at higher temperatures. Dense gas also corresponds to regions where the mass fraction of small grains $f_{\rm S}$ is lower and the mass fraction of carbonaceous grains $f_{\rm C}$ is higher compared to the more diffuse ISM.
@@ 4.2 Dust-to-metal ratio
The evolution over time of the DTM for the fiducial G10LG galaxy is shown in Fig. 3. This quantity is computed for all the gas and dust contained in a cylinder of $r=4\,\rm kpc$ cylindrical radius centred on the box and $h/2=200\,\rm pc$ above and below the disc plane. The DTM rises steeply in less than $50\,\rm Myr$ to its approximately steady-state value at nearly ${\rm DTM}=0.4$ comparable to the canonical ratio in present-day galaxies with solar-like metallicities (e.g. Rémy-Ruyer et al. 2014; De Vis et al. 2019) or in the Milky Way (e.g. Jenkins 2009). The DTM is decomposed into the four simulated dust bins, i.e. into small (dotted) and large (dashed) grains, and into carbonaceous (blue) and silicate (red) grains. The bulk of the DTM is from large grains ($80\,\%$) and from silicates ($66\,\%$), and in particular from large silicate grains ($53\,\%$).
@@ 4.3 Dust size distribution
The top panel of Fig. 5 shows the reconstructed size distribution of grains (in black solid line) from our four populations with size $a_{j}$ ($j$ standing for S or L) and chemical composition $i$ (for carbonaceous or silicate grains), assuming that each population is sampled with a modified log-normal PDF (as in Hirashita 2015 and Hou et al. 2017) of
@@
where $\sigma_{i,j}=0.75$ is the standard deviation of the log-normal part and $C_{i,j}$ is the normalisation of the distribution. The normalisation is given by:
@@
where $\mu=1.22$ is the mean molecular weight of the gas (assuming, for simplicity, that the dusty gas mainly contributing to the extinction curve is fully neutral). The size distribution of grains measured by $\mathcal{D}_{i,j}=\sum_{\Omega}D_{i,j}/\sum_{\Omega}M_{\rm g}$, over a control volume $\Omega$ that is a cylinder of $r=4\,\rm kpc$ radius and $h/2=0.2\,\rm kpc$ semi-height around the centre of the galaxy, is in good agreement with the expected MRN distribution of grains in the Milky Way, also shown in Fig. 5. It shows that the size distribution of carbonaceous and silicate grains are extremely similar, and that the amount of carbonaceous grains is lower than that of silicate grains (as expected from the decomposition of DTM in Fig. 3).
@@ 4.4 Extinction curve
where $N_{\rm H}$ is the hydrogen column number density and $Q_{{\rm ext},i}(a,\lambda)$ is the extinction coefficient that is a function of grain size $a$, wavelength $\lambda$ and grain composition $i$. The size distribution is obtained from the log-normal distribution [^7] We note that the extinction curve is not sensitive to the exact spread of the log-normal reconstruction for the small bin size, but is very sensitive to the one for the large bin (and in particular to that of carbonaceous grains), which should have a sharp cut-off above $a=0.1\,\mu\rm m$ (Weingartner & Draine 2001a). of equation (21). $Q_{{\rm ext},i}(a,\lambda)$ is obtained by using the Mie theory (Bohren & Huffman 1983) with the optical constants for carbonaceous and silicate grains from Weingartner & Draine (2001a). Bottom panel of Fig. 5 shows the extinction curve $A_{\lambda}$ normalised by the extinction $A_{\rm V}$ in the V-band at $\lambda_{\rm V}=0.55\,\rm\mu m$, obtained for the fiducial G10LG simulation at time $t=100$, 200, 300, and $400\,\rm Myr$ for all the dust in the same region (within a cylinder of $r=4\,\rm kpc$ and $h/2=0.2\,\rm kpc$) as used previously for measuring the dust size distribution (top panel of Fig. 5). We also show the scatter of the extinction curve at $t=400\,\rm Myr$ by sampling the corresponding extinction curves for each individual cells within the same region. The extinction curve from the Milky Way analogue G10LG shows an excellent agreement with that observed for the Milky Way (Pei 1992): it shows a similar slope at short wavelengths and the characteristic bump feature at $\lambda=2175\,\rm\AA$ with similar strength. This shape is clearly distinct from the extinction curves from Magellanic Clouds. As expected, the $2175\,\rm\AA$ bump is obtained from the small carbonaceous grains (blue dotted line) due to their optical properties, while the UV-to-optical slope is the result of the small silicate grains (red dotted line). It follows immediately that to obtain steeper slopes and a shallower $2175\,\rm\AA$ bump (as in the Magellanic Clouds extinction curves), the fraction of small carbonaceous grains must be reduced and the fraction of small silicate grains must be enhanced. This will be investigated further in section 5.
@@
Diversity in the extinction curves is also a significant feature of Milky Way extinction curves with typical relative variations across lines-of-sights of $A_{\lambda}/A_{\rm V}$ of $\sim 20\,\%$ at $\lambda^{-1}=8\,\mu\rm m^{-1}$ (Cardelli et al. 1989; Fitzpatrick & Massa 2007), as can be seen in Fig. 5. Although, extreme care must be taken when comparing directly with observational data (line-of-sight effects), we investigate what are the different extinction curves as a function of the various gas phases of the ISM to get a better indication of what are variations of the dust properties across them. In Fig. 6, we show the extinction curves at $t=400\,\rm Myr$ in G10LG, as relative variations of $\delta(A_{\lambda}/A_{\rm V})|_{n}=(A_{\lambda}/A_{\rm V})|_{n}/(A_{\lambda}/A_{\rm V})-1$ at a given $n$ with respect to the global extinction curve $A_{\lambda}/A_{\rm V}$ (as shown in Fig. 5), for different mean gas densities $n\sim 10^{-2},10^{-1},1,10$, and $10^{2}\,\rm H\,cm^{-3}$ (sampled in narrow bins of ±$\pm 0.1\,\rm dex$ around the mean) collected within the same cylinder of $r=4\,\rm kpc$ and $h/2=0.2\,\rm kpc$. Gas at densities of $n\sim 10$ and $10^{2}\,\rm H\,cm^{-3}$ exhibits extinction curves close to the global value. The variation of extinction curves is not monotonic with gas density. The extremely low value of gas density $n\sim 10^{-2}\,\rm H\,cm^{-3}$ corresponds to the larger decrement in the extinction curves: lower bump strength and shallower UV-to-optical slope. This is understood as the consequence of small grain destruction by SN explosions and thermal sputtering. At intermediate densities $n\sim 10^{-1}$ and $1\,\rm H\,cm^{-3}$, the extinction curves approach their global values and stand above the global expectation at $1\,\rm H\,cm^{-3}$ as a signature of shattering. Shattering transfers mass from large grains to small grains preferentially in the diffuse ISM, reinforcing both the bump strength and the UV slope. The highest gas densities are the closest to the global value as this is where most of the dust mass is concentrated ($50\,\%$ of the total dust mass at $n\sim 10\,\rm H\,cm^{-3}$, and $30\,\%$ at $n\sim 10^{2}\,\rm H\,cm^{-3}$). They have lower extinction curves relative to the lower ISM gas density of $n\sim 1\,\rm H\,cm^{-3}$ due to the efficient coagulation of small grains into large grains. Finally the different sampled densities all have large scatter of the order of the mean values, which is indicative of the intrinsic diversity of extinction curves at a given gas density.
Figure 7: Depletion factors as a function of gas density for the G10LG simulation at $400\,\rm Myr$ (black diamonds with standard deviation) compared to the results from Jenkins (2009) (black dashed lines) and from De Cia et al. (2016) (red dashed lines). The solid lines are the rescaled observational fits following Zhukovska et al. (2016). The purple and orange diamonds are the depletion factors for the G10LG simulation using a pyroxene compound (MgFeSi2O6), and an iron-poor olivine compound (Mg1.5Fe0.5SiO4), respectively, instead of the default olivine compound (MgFeSiO4) for silicates. The blue diamonds stand for the G10LG simulation without CO formation. For C depletion, we also show the data points from Jenkins (2009) (triangles) and  Parvathi et al. (2012) (squares).
@@ 4.5 Depletion factors
The presence of dust in the ISM can be characterised by the depletion of the corresponding atomic elements in the gas phase $f_{\rm dep}(X)=1-M_{\rm D}(X)/M_{\rm Z}(X)$, where $M_{\rm D}(X)$ and $M_{\rm Z}(X)$ are respectively the dust mass and the total (dust+gas) mass of the element $X$. We show in Fig. 7 the depletion factor of the various elements as a function of the neutral gas density $n_{\rm neutral}$ (assuming that gas with temperature below $T\leq 10^{4}\,\rm K$ is neutral and fully ionised otherwise) compared to observations. The observational relations are shown with the raw value of Jenkins (2009) as the black dashed lines, and with the recommended renormalisation of their data in black solid line by Zhukovska et al. 2016 (a correction also adopted in Choban et al. 2022) that compensates for the fact that the observed densities are under-estimated since they are averaged along the line-of-sight. We also show the fit provided by Richings et al. (2022) of the data from De Cia et al. (2016), which extends the Jenkins (2009) data to low density lines-of-sight (with or without the rescaling in solid and dashed red lines respectively). For C depletion, we followed Choban et al. (2022), and we reduced the depletion values of Jenkins (2009) by a factor 2 (in that case, we did not rescale in density because the relation is sufficiently shallow) to account for the apparent over-estimate of the C gas-phase abundance from weak compared to strong CII lines (Sofia et al. 2011; Parvathi et al. 2012). We also show the data points from Jenkins (2009) (rescaled by a factor of 2) and the data points from Parvathi et al. (2012) to underline the large uncertainties associated to the estimate of the depletion of C.
As expected, there is more depletion of individual elements in the gas phase ($f_{\rm dep}$ decreases) with increasing density as a result of larger accretion in the dense gas and of destroyed grains in the diffuse phase. Given the large uncertainties in the reconstruction of the observed gas densities, our depletion values are in broad agreement with the data, except for the Mg which clearly stands out of the data, while the other compounds of silicate grains (Fe, Si, and O) are in the observational ballpark. Iron is the most depleted of the four elements entering the composition of silicates, and, thus, is the limiting element in the accretion for the formation of olivine with that type of iron inclusions (i.e. MgFeSiO4). Depletion factors of the silicate-bearing elements are naturally extremely sensitive to the exact composition of the silicate. We performed additional simulations where we have assumed a different silicate compound (MgFeSi2O6 or Mg1.5Fe0.5SiO4). In the pyroxene compound MgFeSi2O6, the mass fraction of Si is $50\,\%$ larger than in the olivine and Fe is $35\,\%$ lower, and, thus, Si becomes the limiting element. Indeed, Fig. 7 shows that $f_{\rm dep}({\rm Si})$ is lower for a G10LG simulation that is performed assuming pyroxene compound instead of olivine, and that Si, then, becomes the lowest of the silicate-bearing elements. Similarly, decreasing the amount of iron inclusions in olivine (or pyroxene) changes the picture: a G10LG simulation assuming iron-poor olivine Mg1.5Fe0.5SiO4 instead of iron-rich olivine MgFeSiO4 leads to lower values of $f_{\rm dep}({\rm Mg})$ that makes Mg the new limiting element of silicate grains.
@@
For C depletion, our resulting values are well within the range of observations although the observational data exhibits a large scatter. We also highlight the role of CO formation at high gas densities in limiting the depletion of C elements by grains (see also Choban et al. 2022). If we do not account for CO formation in the limitation of carbonaceous dust growth from ISM accretion, i.e. as for silicate grains, carbonaceous grain growth is only limited by ice mantle coating starting above $n=10^{4}\,\rm H\,cm^{-3}$ (G10LG_NCO simulation), then carbonaceous grains accrete much more efficiently from the ISM, leading to too much depletion of C. Therefore, CO formation indirectly allows for more moderate values of $f_{\rm dep}({\rm C})$ in better agreement with the data.
@@ 4.6 Which process dominates the dust growth?
Figure 9: Dust-to-metal ratio (DTM, top left), dust-to-gas ratio (DTG, top right), fraction of small grains $f_{\rm S}$ (bottom left), and fraction of carbonaceous grains $f_{\rm C}$ (bottom right) as a function of density $n$ and temperature $T$ for the G10LG galaxy at time $t=400\,\rm Myr$. Black contours are for the mass distribution of gas.
Figure 10: Top panel: dust-to-metal ratio (solid lines, left axis) and gas metallicity (dashed lines, right axis) as a function of time for the 4 different initial metallicities of the G10LG galaxy: $0.1$, $0.3$, $1$, and $2\,\rm Z_{\odot}$ (from bottom to top). Bottom panel: fraction of carbonaceous grains (solid lines) and of small grains (dot-dashed lines). The DTM increases with metallicity, the fraction of carbonaceous grains increases with metallicity and the fraction of small grains decreases with metallicity.
Figure 11: Comparison of extinction curves in the G10LG galaxy at time $t=400\,\rm Myr$ for different initial metallicities ($Z_{\rm g,0}=0.1$, $0.3$, $1$, and $2\,\rm Z_{\odot}$ for G10LG_VLZ, G10LG_LZ, G10LG, and G10LG_HZ respectively). The top panel shows the extinction curves and the bottom panel shows the relative variation of the extinction curve with respect to the fiducial simulation G10LG. Galaxies with lower metallicities produce steeper UV-to-optical slopes due lower accretion efficiencies in the dense gas where coagulation proceeds, hence, the fraction of small grains is larger.
@@
Figure 13: Comparison of extinction curves in the G10LG galaxy at time $t=400\,\rm Myr$ for variations in coagulation (red), shattering rates (blue), and in SN dust destruction efficiency for small grains (cyan). An increase (decrease) in the corresponding rate is depicted by a dashed (respectively solid) line. The triple-dot-dashed orange line stands for an increased in both the shattering and the coagulation rates. Top panel shows the extinction curves and the bottom panel shows the relative variation of the extinction curve with respect to the fiducial simulation.
@@
The evolution of the dust mass in the simulated galaxies obeys different growth and destruction processes. Figure 8 shows the different dust mass variation rates for different processes involved in the dust evolution of the fiducial G10LG run: accretion, thermal sputtering, SN shock destruction, dust released by stars through stellar winds in SN ejecta, astration, coagulation and shattering. All rates are obtained within the whole simulation box, and curves are smoothed over a 5 Myr timescale for sake of readability. Apart from the first initial peak at a few $10\,\rm Myr$ after the start of the simulation, all curves show little variation with time due to the low SFR and no significant evolution in gas-phase metallicity (which could affect the accretion time since $t_{\rm acc}\propto Z^{-1}$). Accretion of the available refractory elements from the ISM onto the small grains is by far the dominant process of dust mass growth compared to dust released in the ejecta of SNe or of stellar winds (as expected for sufficiently metal-rich galaxies, e.g. Popping et al. 2017; Aoyama et al. 2017; Vijayan et al. 2019; Choban et al. 2022), or by accretion onto large grains. Destruction by SNe is a factor of 3 below the mass growth by accretion and is the dominant mechanism for dust mass removal. Thermal sputtering, which is the mechanism by which the dust is destroyed by pockets of hot ($T>10^{6}\,\rm K$) and diffuse ($n<0.01\,\rm H\,cm^{-3}$) gas, is a factor 2 below the destruction rate by SNe. Finally destruction of the dust by astration is lower by a factor 4 compared to destruction by SNe. Interestingly, the thermal sputtering removes more mass on large grains than on small grains, and SNII destruction removes an equal mass of small and large grains, while their physical timescale for dust destruction is an increasing function of the grain size. Hence, one would naively expect the sputtering to remove more dust mass from small grains. However, since the dust mass is 4 times larger for large grains than small grains (see Fig. 3), this effect is compensated by the available reservoir of dust mass in each bin of grain size.
@@
4.7 Dust grain distribution in the multiphase interstellar medium